# “It Felt Like I Was Left in the Dark”: Exploring Information Needs and Design Opportunities for Family Caregivers of Older Adult Patients in Critical Care Settings

### v2

@@ 7.3. Risks and Ethical Concerns of AI-based Clinical Information Technologies
A central tension emerges from our prototype’s dual role as both a support tool and a potential stressor. Caregivers acknowledged that our prototype could facilitate their access to information, such as tracking changes in a patient’s condition, yet they also expressed concerns about information overload, particularly when frequent, non-essential updates increased their cognitive burden. This aligns with previous research on AI-supported decision tools, which found that frequent notifications and complex data presentations can inadvertently increase user stress (Wang et al., 2023). While AI can automate data collection and generate insights, human oversight remains crucial to interpret alerts, prioritize key information, and prevent caregivers from becoming overwhelmed.
@@ 7.4. Limitations and Future Work
Our study has several limitations that should be acknowledged. First, the scope of our participant pool was limited, with 15 participants in the formative study and 10 participants in the evaluation phase. While these participants provided valuable insights, the small sample size may limit the generalizability of our findings. Future research should aim to include a larger and more diverse group of caregivers to ensure broader applicability. Nevertheless, our analysis reached a point of saturation, ensuring that the results comprehensively reflect the experiences and perspectives of caregivers for older adult patients in critical care settings. Second, the prototype developed in this study is a proof-of-concept and has not been integrated with Electronic Health Record (EHR) systems or deployed in real-world clinical settings. While the prototype demonstrates the potential of AI-based tools to support caregivers, its effectiveness and usability in practical, high-stakes environments remain untested. Future work should focus on integrating such tools with existing healthcare infrastructures and evaluating their performance in real-world scenarios.
Additionally, our study specifically focused on designing a prototype to address caregiver-facing challenges. While this approach provides valuable insights into the needs of family caregivers, it does not fully account for the perspectives of clinicians or other stakeholders in the care process. Future studies should involve clinicians to explore how such tools can facilitate collaboration between caregivers and healthcare providers, ensuring that the design meets the needs of all parties involved. Lastly, while our research primarily addresses the ICU setting, there is a broader need to explore how similar tools can be adapted for other clinical environments. The challenges faced by caregivers in ICUs may differ from those in other settings, such as long-term care facilities or home-based care. Future work should delve deeper into these contexts to develop more versatile solutions that can support caregivers across a wide range of clinical scenarios.
@@ 8. Conclusion
In this study, we investigated the information needs of family caregivers of ICU older adult patients and designed an AI-based prototype, CareMap, to address their challenges in accessing and understanding clinical information. Through a formative study with 15 caregivers, we first uncovered the multifaceted challenges caregivers face in the ICU. Building on these insights, we designed an AI-based prototype and evaluated it with 10 caregivers. Results from the evaluation study suggested that, despite participants identifying potential trust risks in using an AI-based system, they emphasized that the plain-language explanations and unified visual timeline empowered them to more effectively access and understand information in critical care settings. This paper contributes to HCI research in healthcare by highlighting the need for AI systems that actively engage caregivers and ensure they feel informed and supported throughout the care process, rather than sidelined.
@@ Appendix A Formative Study - Interview Script
Appendix B Evaluation Study - Interview Script
The following set of questions was asked for each interface presented to the participants.
  • Question 1 - Feedback on the Design: (After showing a screenshot of the UI) How do you like this interface?
  • Question 2 - Utility: Do you think this function would be useful to you? Why or why not?
  • Question 3 - Use Cases: How would you use this interface/function? [What kind of questions would you ask this chatbot?]
  • Question 4 - Improvement: What improvements can be made?

### v3

@@ 7.3. Risks, Technical and Ethical Concerns of AI-based Clinical Information Technologies
7.3.1. Infrastructure and Technical Challenges
Successful real-time implementation of CareMap in clinical environments requires coordinated efforts across both infrastructure and healthcare teams. From the perspective of infrastructure, many mHealth applications—such as MyChart (Ramsey et al., 2018)—have already been integrated into standard clinical workflows and tightly coupled with mainstream electronic health record (EHR) systems. This trend demonstrates that healthcare institutions possess mature data interfaces, robust security protocols, and established mechanisms for workflow integration. As a result, incorporating the conversational capabilities of CareMap into existing EHRs or remote monitoring platforms is both technically and institutionally viable.
Equally important is the sustained engagement of healthcare teams to ensure the quality and reliability of information, as well as the overall stability of system operations. AI-based clinical technologies can support the interpretation of medical information and the generation of initial responses, but in the high-risk, continuously monitored environment of the ICU, human oversight remains essential to preserve medical accuracy and contextual relevance. Clinical involvement also enhances information credibility and allows for critical judgment in abnormal or emergency situations. In parallel, system deployment must address ethical concerns and workflow integration challenges, including the need to prevent additional cognitive or operational burdens for healthcare team (Wu et al., 2024; Zhang et al., 2024).
7.3.2. Ethical Concerns of AI-based Clinical Information Technologies
Another central tension emerges from our prototype’s dual role as both a support tool and a potential stressor. Caregivers acknowledged that our prototype could facilitate their access to information, such as tracking changes in a patient’s condition, yet they also expressed concerns about information overload, particularly when frequent, non-essential updates increased their cognitive burden. This aligns with previous research on AI-supported decision tools, which found that frequent notifications and complex data presentations can inadvertently increase user stress (Wang et al., 2023).
@@ 7.4. Limitations and Future Work
Our study has several limitations that should be acknowledged. First, our participant pool was relatively limited, consisting of 15 family caregivers in the formative study and 10 family caregivers and 2 clinicians in the evaluation phase. While these participants provided valuable insights, the small sample size may limit the generalizability of our findings. Future research should aim to include a larger and more diverse group of caregivers and clinicians to ensure broader applicability. Nevertheless, our analysis reached a point of saturation, and participants contributed rich and diverse perspectives that were sufficient for the exploratory and proof-of-concept stage of this research. This ensures our study results comprehensively reflect the experiences and perspectives of caregivers for older adult patients in critical care settings.
Second, the prototype developed in this study is a proof-of-concept and has not been integrated with Electronic Health Record (EHR) systems or deployed in real-world clinical settings. While the prototype demonstrates the potential of AI-based tools to support caregivers as discussed in Sec 7.3.1, its effectiveness and usability in practical, high-stakes environments remain untested. Future work should focus on integrating such tools with existing healthcare infrastructures and evaluating their performance in real-world scenarios.
Lastly, while our research primarily addresses the ICU setting, there is a broader need to explore how similar tools can be adapted for other clinical environments. The challenges faced by caregivers in ICUs may differ from those in other settings, such as long-term care facilities or home-based care. Future work should delve deeper into these contexts to develop more versatile solutions that can support caregivers across a wide range of clinical scenarios.
@@ 8. Conclusion
In this study, we investigated the information needs of family caregivers of ICU older adult patients and designed an AI-based prototype, CareMap, to address their challenges in accessing and understanding clinical information. Through a formative study with 15 caregivers, we first uncovered the multifaceted challenges caregivers face in the ICU. Building on these insights, we designed an AI-based prototype and evaluated it with 10 caregivers and 2 clinicians. Results from the evaluation study suggested that, despite participants identifying potential trust risks in using an AI-based system, they emphasized that the plain-language explanations and unified visual timeline empowered them to more effectively access and understand information in critical care settings. This paper contributes to HCI research in healthcare by highlighting the need for AI systems that actively engage caregivers and ensure they feel informed and supported throughout the care process, rather than sidelined.
@@ Appendix A Formative Study - Interview Script
Appendix B Evaluation Study
B.1. Interview Script
The following sets of questions were asked for each interface presented to the participants.
Family Caregivers
  • Question 1 – Feedback on the Design: (After showing a screenshot of the UI) How do you like this interface?
  • Question 2 – Utility: Do you think this function would be useful to you? Why or why not?
  • Question 3 – Use Cases: How would you use this interface/function? [What kind of questions would you ask this chatbot?]
  • Question 4 – Improvement: What improvements can be made?
Clinical Experts (Physicians)
  • Question 1 – Output Quality: How accurate, clear, and clinically appropriate do you find the system-generated outputs (daily summaries, treatment plans, chatbot responses)?
  • Question 2 – Clinical Utility: Do you think this system could support clinicians or improve communication with caregivers? Why or why not?
  • Question 3 – Usability Feedback: From a clinical perspective, how usable and trustworthy do you find this interface?
  • Question 4 – Improvement: What improvements would you suggest to make the system more clinically reliable and usable?
B.2. Questionnaire - Clinician Feedback on AI-generated
Please rate your level of agreement with the following statements about the AI-generated summary. 1 - strongly disagree, 5 - strongly agree.
  • Overall Accuracy - The AI-generated output presents medically accurate information without major errors.
  • Hallucination - The AI-generated output does not contain fabricated or factually incorrect information.
  • Readability - The AI-generated output is written in clear, well-structured language that is easy to read and understand quickly.
  • Specificity - The AI-generated output provides sufficient specific details (e.g., vital signs, medication names, test results) rather than vague or general statements.
  • Overall Quality - Overall, the AI-generated output is clear, trustworthy, and helpful for quickly understanding the patient’s daily status.